Please help with C answers provided for A and B

The ledger of American Company has the following work in process account.

Work in Process—Painting

5/1 Balance 3,790 5/31 Transferred out ?
5/31 Materials 5,700
5/31 Labor 3,156
5/31 Overhead 2,080
5/31 Balance ?


Production records show that there were 600 units in the beginning inventory, 30% complete, 1,600 units started, and 1,500 units transferred out. The beginning work in process had materials cost of $2,440 and conversion costs of $1,350. The units in ending inventory were 40% complete. Materials are entered at the beginning of the painting process.

(a)

Correct answer iconYour answer is correct.

How many units are in process at May 31?

Work in process, May 31

enter a number of units in process at May 31

units

700

eTextbook and Media

  

Attempts: 1 of 5 used

(b)

Correct answer iconYour answer is correct.

What is the unit materials cost for May? (Round unit cost to 2 decimal places, e.g. 2.25.)

The unit materials cost for May

$enter a dollar amount of the unit materials cost for May rounded to 2 decimal places

3.7

eTextbook and Media

  

Attempts: 1 of 5 used

(c)

What is the unit conversion cost for May? (Round unit cost to 2 decimal places, e.g. 2.25.)

The unit conversion cost for May

$enter a dollar amount of the unit conversion cost for May rounded to 2 decimal places

?
Answer #1

1) Ending WIP = 600+1600-1500 = 700 Units

2) Equivalent unit of material = 1500+700*100% = 2200

Unit cost of material = (2440+5700)/2200 = 3.7

3) Equivalent unit of conversion = 1500+(700*40%) = 1780

Unit cost of conversion = (1350+3156+2080)/1780 = 3.7